OpenStudy (anonymous):

Of teens 225 or 9/20 said they prefer chocolate. About how many teens surveyed?

OpenStudy (anonymous):

let x be the number of teens surveyed 9x/20 = 225 find x

TheSmartOne (thesmartone):

\[\frac{ 9 }{ 20 }x=225\] \[x=225 \times \frac{ 20 }{ 9 }\] \(\ x=?\)

TheSmartOne (thesmartone):

225/9=25 So \[x=25 \times 20=500\]